OpenStudy (anonymous):

EASY FAN/MEDAL Solve for x: |x + 2| + 16 = 14 A)x = −32 and x = −4 B)x = −4 and x = 0 C)x = 0 and x = 28 D)No solution

OpenStudy (anonymous):

(x+2)+16=14 x+2=-2 x+2=-2 x=-4 x=0 snswer is B

OpenStudy (anonymous):

THANKS ^-^

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can I get a medal? :D

myininaya (myininaya):

\[|x+2|+16=14 \\ \text{ subtract 16 on both sides } \\ |x+2|=14-16 \\ |x+2|=-2 \] but absolute value don't spit out negative numbers

myininaya (myininaya):

so what do you think the solution is or do you think there is none @shawty247

myininaya (myininaya):

I did just say absolute value function doesn't spit out negative numbers